source voltage= 40V

series resistance = 200 ohm

zener voltage= 15V


1
Expert's answer
2022-02-09T16:18:57-0500




"V_{In}=40V," "R_s=200\Omega"

"V_z=15V"

and let "R_L=500\Omega"


"V_{across} \>R_s=40-15"

"=25V"


"I_s=\frac{25}{200}=125mA"


"I_L=\frac{15}{500}=30mA"


"\implies\>I_z=125-30mA"

"=95mA"